Market Street Sweets sells two types of Pralines:a carmel version for 1.65 per pound and a chocolate version for 1.25 per pound. A Mixture of these pralines can be sold for 1.49 per pound. How many pounds of each kind should be used to make 5 pounds of the mixed pralines?
Thank you

Answer by palanisamy(496)

Let x pounds of Pralines:a carmel version and y pounds of chocolate version be mixed
Then total weight is x+y=5 ...(1)
Next, price is 1.65x+1.25y = (1.49)*5
165x+125y=745 ...(2)
Now, (1)*125=> 125x+125y = 625 ..(3)
(2)-(3)=> 40x = 120
x=120/40=3
y=5-3=2
So, 3 pounds of Pralines:a carmel version and 2 pounds of chocolate version are mixed